OpenStudy (anonymous):

.Choose the equation of the line passing through the point (-3, 1) and parallel to y = -x - 4. y = x + 2 y = -x - 2 y = -x + 4 y = x - 4

OpenStudy (anonymous):

Firstly find the slope of the parallel line given..

OpenStudy (anonymous):

y = -x - 4 Compare it with y = mx + c And tell me what is m here ??

OpenStudy (anonymous):

1º it will have to have the same slope: only answers left are: y = -x - 2 and y = -x + 4 2º check if the point (-3, 1) satisfies the equations above.

OpenStudy (anonymous):

i just dont get how to solve it at all :( im horrible at math

OpenStudy (anonymous):

This might be easier for you if you use point-slope form instead. \[y - y_1 = m(x - x_1)\]Remember that slopes of parallel lines are the same. So the slope here is -1. The, remember that \((x_1, y_1)\). So \(x_1 = -3\) and that \(y_1 = 1\). \[y - 1 = -1(x - 3)\]\[y = -1(x - 3) + 1\]Can you simplify the right side?

OpenStudy (anonymous):

@Calcmathlete there is little mistake I guess..

OpenStudy (anonymous):

Oh ok. \[y - 1 = -1(x + 3)\]\[y = -1(x + 3) + 1\]Can you simplify this?